OpenStudy (anonymous):

find the slope if it exists of the line containing the pair of points. (5,0), (0,-1)

OpenStudy (anonymous):

.3/6

OpenStudy (mertsj):

\[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (mertsj):

That is the slope formula. Plug in your points and simplify.

OpenStudy (anonymous):

3/6?is that it?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

Yes.

OpenStudy (mertsj):

Subtract the y values. What is 0 - (-1)

OpenStudy (anonymous):

THE ANSWER IS 3/6!!!!

OpenStudy (abmon98):

-1-0/0-5=1/5 thats is the gradient

OpenStudy (anonymous):

0

OpenStudy (mertsj):

0-(-1)=0+1=1 So when you subtract the y values, you get 1 Now subtract the x values. What is 5-0?

OpenStudy (anonymous):

5

OpenStudy (mertsj):

Put 1 over 5. That is the slope. what is your answer?

OpenStudy (anonymous):

1/5

OpenStudy (mertsj):

Yes. Good. Now tell the answer to @madln45

OpenStudy (anonymous):

lol you can tell them

OpenStudy (mertsj):

You need to practice this concept a little bit more. It is important that you learn how to find the slope when you know two points.